# Trees of Graphs as Boundaries of Hyperbolic Groups

### v1

@@ 7. $V$-trees of graphs and their decompositions into core and arms
In this section we begin preparations for the proof of the implication (3$\Rightarrow$(2) in Theorem A (and the implication (3$\Rightarrow$(2) of Theorem B). The V-tree systems discussed in this section should be thought of as subsystems in the regular tree systems of graphs, and their closer understanding is also a first step in understanding the latter, and in relating the latter with Gromov boundaries of the considered hyperbolic groups.
@@
    $\Gamma$ is a finite graph which is essential (i.e. has no isolated vertex), and we assume it has no loop edges;
@@
We now pass to describing certain tree system $\Theta[{\mathcal{V}}]=\Theta[\Gamma,{\mathbf{a}}]$ canonically associated to a connecting $V$-system $\mathcal{V}$. Its limit $\lim\Theta[{\mathcal{V}}]$ will be called the $V$-tree of graphs for $\mathcal{V}$, and it will be denoted ${\mathcal{X}}_{V}({\mathcal{V}})$, or ${\mathcal{X}}_{V}(\Gamma,{\mathbf{a}})$, where the role of the subscript $V$ is just to indicate the $V$-tree nature of the resulting space. Actually, the $V$-tree ${\mathcal{X}}_{V}({\mathcal{V}})$ will be natuarally a space with peripherals, and the description of a family of peripherals will be a part of its definition.
Similarly as in Section 5, we consider a standard family ${\mathcal{N}}_{\Gamma}$ of normal neighbourhoods in $\Gamma$, as described in Example 2.2, and the space with peripherals $\Gamma^{\circ}$, called the punctured graph $\Gamma$, described in the same example. Consider also the set $\Omega^{V}_{\Gamma}$ of the so called $V$-peripherals of $\Gamma^{\circ}$, as described right after Definition 5.1.
@@
    for each black vertex $t$ of $T$ the constituent space with peripherals at $t$ is equal to the pair $(\Gamma^{\circ},\Omega^{V}_{\Gamma})$; this means in particular that the vertices in $N_{t}$ (i.e. the vertices adjacent to $t$) are in the fixed bijective correspondence with the family $\Omega_{\Gamma}^{V}$;
@@
    for each white vertex $u$ of $T$, denote by $e,e^{\prime}$ the two edges of $T$ adjacent to $u$, and by $P_{e},P_{e^{\prime}}$ the $V$-peripherals of $\Gamma_{b(e)}^{\circ}$ and $\Gamma_{b(e^{\prime})}^{\circ}$ associated to $u$ via the bijective correspondences mentioned above in condition (v1); then the set $\Sigma_{u}$ and the bijections $\varphi_{e},\varphi_{e^{\prime}}$ are given so that the composition $\varphi_{e^{\prime}}\varphi_{e}^{-1}:P_{e}\to P_{e^{\prime}}$ coincides with the map $\alpha_{v}$, where $v$ and $a(v)$ are the vertices of $\Gamma$ corresponding to the $V$-peripherals $P_{e},P_{e^{\prime}}$, respectively, and where we refer to the natural identifications $P_{e}={\rm Lk}(v,\Gamma)$ and $P_{e^{\prime}}={\rm Lk}(\alpha(v),\Gamma)$.
We skip a straightforward justification of the fact that a tree system satisfying the above requirements exists and is unique up to an isomorphism of tree systems (the argument is similar to that presented in Section 6 for the tree system $\Theta[{\mathcal{R}}]$).
Now, let ${\mathcal{X}}_{V}({\mathcal{V}}):=\lim\Theta[{\mathcal{V}}]$. We describe the family $\Omega_{V}({\mathcal{V}})$ of natural peripheral subsets for ${\mathcal{X}}_{V}({\mathcal{V}})$. Let $P$ be any $E$-peripheral in any constituent space $\Gamma^{\circ}_{t}$. Clearly, $P$ is then naturally a subset of the limit $\lim\Theta[{\mathcal{V}}]$. Moreover, for any two distinct $E$-peripherals $P,P^{\prime}$ as above, the induced subsets of $\lim\Theta[{\mathcal{V}}]$ are disjoint. It is also not hard to realize (and we omit the details) that the family $\Omega_{V}({\mathcal{V}})$ of all such subsets of $\lim\Theta[{\mathcal{V}}]$ is null, and thus we take it as the family of natural peripherals for ${\mathcal{X}}_{V}({\mathcal{V}})$.
@@
A graph which coincides with a single edge will be called the segment, and we donote this graph by $I$. The punctured segment $I^{\circ}$ is then the space with peripherals in which two of the peripheral sets are singletons, while the others are doubletons. If we delete from the set of peripherals of $I^{\circ}$ the two singleton peripherals, we get a new space with peripherals (with the same underlying space, but smaller family of peripherals) which we call the internally punctured segment, and which we denote $I^{\circ}_{\rm int}$. Alternatively, $I^{\circ}_{\rm int}$ can be described as obtained by deleting from $I$ the interiors of a dense and null family of pairwise disjoint normal neighbourhoods of points lying in the interior of $I$, and by taking the boundaries of those neighbourhoods as peripherals.
Observe that there are exactly two essentially distinct connecting $V$-systems ${\mathcal{V}}=(\Gamma,{\mathbf{a}})$ with $\Gamma=I$: the first one with the involution $a:V_{I}\to V_{I}$ equal to the identity of the vertex set $V_{I}$, and the second one with $a$ equal to the transposition of the doubleton $V_{I}$. In both cases the resulting $V$-tree ${\mathcal{X}}_{V}({\mathcal{V}})$ will be called $V$-tree of segments. Moreover, we obviously have the following.
@@
We now pass to describing some natural decomposition of any $V$-tree of graphs ${\mathcal{X}}_{V}({\mathcal{V}})$ into convenient pieces. These pieces will be called the core and the arms of ${\mathcal{X}}_{V}({\mathcal{V}})$, and they will have a form of spaces with peripherals. We will show (in Lemma 7.12) that decomposition into these pieces has tree-like nature, i.e. that the space ${\mathcal{X}}_{V}({\mathcal{V}})$ is canonically homeomorphic to certain tree of these pieces. This new expression of ${\mathcal{X}}_{V}({\mathcal{V}})$ as tree of spaces does not coincide with the original expression of this space as $V$-tree of graphs, and this reinterpretation of the spaces of form ${\mathcal{X}}_{V}({\mathcal{V}})$ will be one of our tools in the proof of the implications (3$\Rightarrow$(2) in Theorem A, and (3$\implies$(2) in Theorem B.
Before describing the pieces and the decomposition, we need to introduce more terminology. A punctured edge of a punctured graph $\Gamma^{\circ}$ is the intersection of $\Gamma^{\circ}$ (viewed as subspace of $\Gamma$) with any edge $e$ of $\Gamma$, and we denote such a punctured edge by $e^{\circ}$. Obviously, since we assume that $\Gamma$ contains no loop edges, each $e^{\circ}$ has naturally the structure of a space with peripherals with which it is homeomorphic to the punctured segment $I^{\circ}$. The endpoints of a punctured edge $e^{\circ}$ are the two points of $e^{\circ}$ which correspond to the two singleton peripherals of this space. The set of all endpoints in all punctured edges of $\Gamma^{\circ}$ is clearly in a natural bijective correspondence with the union of all vertex links of the graph $\Gamma$, and we will identify each such endpoint with the corresponding point in the appropriate vertex link. Now, given a $V$-tree system $\Theta[{\mathcal{V}}]=(T,\{\Gamma_{t}^{\circ}\},\{\Sigma_{u}\},\{\varphi_{e}\})$, we say that punctured edges $e_{1}^{\circ}$ of $\Gamma_{t_{1}}^{\circ}$ and $e_{2}^{\circ}$ of $\Gamma_{t_{2}}^{\circ}$ are $\Theta$-adjacent if $t_{1},t_{2}$ are distinct black vartices of $T$ with common adjacent white vertex, which we denote $u$, and if for appropriate endpoints $x_{1},x_{2}$ of $e_{1}^{\circ},e_{2}^{\circ}$ (namely those belonging to the images $\varphi_{[t_{1},u]}(\Sigma_{u})$ and $\varphi_{[t_{1},u]}(\Sigma_{u})$) we have $\varphi_{[t_{1},u]}^{-1}(x_{1})=\varphi_{[t_{2},u]}^{-1}(x_{2})$. Obviously, the points $x_{1},x_{2}$ at which the edges $e_{1}^{\circ},e_{2}^{\circ}$ are adjacent get identified in the limit space ${\mathcal{X}}_{V}({\mathcal{V}})$, which explains the name chosen for $\Theta$-adjacency.
@@
A line in a $V$-tree system $\Theta[{\mathcal{V}}]$ is a bi-infinite sequence $(e_{n}^{\circ})_{n\in Z}$ of punctured edges from the punctured graphs $\Gamma_{t}^{\circ}:t\in V^{b}_{T}$ such that for each $n$ the edge $e_{n}^{\circ}$ is adjacent to $e_{n-1}^{\circ}$ at one of its endpoints, and it is adjacent to $e_{n+1}^{\circ}$ at its other endpoint. We do not distinguish sequences as above which differ by a shift or by inversion of the indices (i.e. we view two such sequences as representing the same line).
Obviously, since $\Gamma$ has no loop edges, any punctured edge in any constituent graph $\Gamma_{t}^{\circ}$ belongs to a unique line in $\Theta[{\mathcal{V}}]$. It is also not hard to observe that in each $V$-tree system $\Theta[{\mathcal{V}}]$ for which $\Gamma$ has more than two vertices, there are many (actually countably infinitely many) lines. We denote the set of all lines in $\Theta[{\mathcal{V}}]$ by ${\mathcal{L}}({\mathcal{V}})$. Since each edge $e_{n}^{\circ}$ of any line $L$ is a punctured edge in some copy $\Gamma_{t}$ of the graph $\Gamma$, we will often refer to it as appearance in $L$ of the corresponding (unpunctured) edge of $\Gamma$.
Let $L=(e^{\circ}_{n})_{n\in Z}$ be a line in $\Theta[{\mathcal{V}}]$, and for each $n$ denote by $t_{n}$ this black vertex of $T$ for which $e_{n}^{\circ}$ is a punctured edge of $\Gamma_{t_{n}}^{\circ}$, and by $u_{n}$ the unique white vertex of $T$ adjacent to both $t_{n-1}$ and $t_{n}$. The bi-infinite sequence
@@
is then a (bi-infinite) combinatorial geodesic in $T$, which we denote by $\gamma_{L}$ and call the geodesic induced by $L$. We make a record of the following properties of lines and their induced geodesics, whose straightforward proofs we omit. (These proofs are based on the easily seen observation that lines, viewed as sequences of appearances of edges from $\Gamma$, are periodic.)
@@
An arm in a $V$-tree of graphs ${\mathcal{X}}_{V}({\mathcal{V}})$ is its subspace $A_{L}$ of the following form. Given any line $L\in{\mathcal{L}}({\mathcal{V}})$, $L=(e_{n}^{\circ})_{n\in Z}$, take as $A_{L}$ the union of the set $\partial\gamma_{L}\subset\partial T$ (consisting of the two ends of the geodesic $\gamma_{L}$) and the image in #$\#\Theta[{\mathcal{V}}]$ of the union $\bigcup_{n\in Z}e_{n}^{\circ}$. We view $A_{L}$ as a space with peripherals, where the family of peripherals consists of only one set, namely $\partial\gamma_{L}$. We distinguish also in $A_{L}$ some other subsets, namely the ones corresponding to all doubleton peripherals in all punctured edges $e_{n}^{\circ}$ constituting $A_{L}$; we call these subsets the supplementary peripherals of $A_{L}$, in contrast with the above mentioned subset $\partial\gamma_{L}$, which we call the basic peripheral of $A_{L}$.
For each arm $A_{L}$ we consider the subspace topology induced from ${\mathcal{X}}_{V}({\mathcal{V}})$. It is not hard to observe that, when equipped with this topology, any arm $A_{L}$ naturally gets the form of a $V$-tree of segments. In particular, we obtain the following.
@@
The core of ${\mathcal{X}}_{V}({\mathcal{V}})=\lim\Theta[{\mathcal{V}}]$ is the subspace $C({\mathcal{V}})$ corresponding to the subset $\partial T\subset\lim\Theta[{\mathcal{V}}]$. It is equipped (as part of its structure) with the family of doubleton subsets $\Omega^{c}_{\mathcal{V}}:=\{\partial\gamma_{L}:L\in{\mathcal{L}}({\mathcal{V}})\}$ (with which it is a space with peripherals, see Example 7.11 below).
@@
    The family $\Omega^{c}_{\mathcal{V}}$ of subsets of the core is null. To see this, it is sufficient to note that the distances in $T$ from a fixed base point of the geodesics $\gamma_{L}$ diverge to $\infty$. This however easily follows from the fact that only finitely many geodesics of the form $\gamma_{L}$ passes through any vertex of $T$, and the fact that $T$ is locally finite. The family $\Omega^{c}_{\mathcal{V}}$ consists also of sets that are pairwise disjoint, see Lemma 7.5(1).
@@
We now generalize V-trees of segments to V-trees of $\theta$-graphs. We start with describing the $\theta$-graphs - a family of graphs which appear explicitely in the statement of Theorem B, and which appear also implicitely in Theorem A, since the regular trees of graphs mentioned in conditions (2) and (3) of this result essentially involve certain $\theta$-graphs (compare Theorem 10.1 and the description of the graphical connecting system ${\mathcal{R}}_{G}$ in Section 10).
@@
Generalizing the exposition of Example 7.2, we define the internally punctured $\theta$-graph $(\Xi_{k})^{\circ}_{{\rm int}}$ as the space with peripherals obtained from $\Xi_{k}$ by deleting the interiors of any dense family of pairwise disjoint normal neighbourhoods of non-vertex points of $\Xi_{k}$, and by taking the boundaries of those neighbourhoods as peripherals. Such a space is obviously unique up to a homeomorphism respecting peripherals, its all peripherals are doubletons, and vertices of $\Xi_{k}$ belong to this space, and are not contained in any peripheral.
Observe that any connecting $V$-tree system ${\mathcal{V}}=(\Xi_{k},{\mathbf{a}})$ of $\theta$-graphs has precisely $k$ lines $L_{1},\dots,L_{k}$. Moreover, the arms $A_{1},\dots,A_{k}$ of the corresponding $V$-tree of $\theta$-graphs ${\mathcal{X}}_{V}({\mathcal{V}})$ have common set of endpoints, and this set (a doubleton) coincides with the core $C({\mathcal{V}})$ of this $V$-tree. The set of peripherals of ${\mathcal{X}}_{V}({\mathcal{V}})$ coincides with the union of the sets of supplementary peripherals of the arms $A_{i}$. Thus, as a space with peripherals, ${\mathcal{X}}_{V}({\mathcal{V}})$ is homeomorphic to the internally punctured $\theta$-graph $(\Xi_{k})^{\circ}_{\rm int}$. The core $C({\mathcal{V}})$ of this $V$-tree coincides then with the vertex set of $\Xi_{k}$ (viewed as the subset of $(\Xi_{k})^{\circ}_{\rm int}$). The only peripheral of this core coincides with the whole of the core.
@@
As the constituent spaces for $\Psi[{\mathcal{V}}]$ we take $K_{t_{c}}=C({\mathcal{V}})$ and $K_{t_{A}}=A$. As the peripheral spaces for $\Psi[{\mathcal{V}}]$ we take $\Sigma_{u_{P}}=P$, and as the maps $\varphi_{[u,t]}:\Sigma_{u}\to K_{t}$ we take the natural inclusions of forms $P\subset C[{\mathcal{V}}]$ and $\partial\gamma_{L}\subset A_{L}$.
@@
The inclusions of form $A_{L}\subset{\mathcal{X}}_{V}({\mathcal{V}})$, for all $L\in{\mathcal{L}}({\mathcal{V}})$, and $C[{\mathcal{V}}]\subset{\mathcal{X}}_{V}({\mathcal{V}})$ induce a well defined map $i_{\mathcal{V}}:\lim\Psi[{\mathcal{V}}]\to{\mathcal{X}}_{V}({\mathcal{V}})$, and this map is a homeomorphism.

### v2

@@ 7. $V$-trees of graphs and their decompositions into core and arms
In this section we begin preparations for the proof of the implication (3$\Rightarrow$(2) in Theorem A (and the implication (3$\Rightarrow$(2) of Theorem B). The V-tree systems as discussed in this section will appear later as subsystems in some special regular tree systems of graphs that we will be considering, and their closer understanding is thus a first step in understanding the latter, and in relating the latter with Gromov boundaries of hyperbolic groups that are under our interest.
@@
    $\Gamma$ is a finite graph which is essential (i.e. is nonempty and has no isolated vertex), and we assume it has no loop edges;
@@
We now pass to describing a certain tree system $\Theta[{\mathcal{V}}]=\Theta[\Gamma,{\mathbf{a}}]$ canonically associated to a connecting $V$-system $\mathcal{V}$. Its limit $\lim\Theta[{\mathcal{V}}]$ will be called the $V$-tree of graphs for $\mathcal{V}$, and it will be denoted ${\mathcal{X}}_{V}({\mathcal{V}})$, or ${\mathcal{X}}_{V}(\Gamma,{\mathbf{a}})$, where the role of the subscript $V$ is just to indicate the $V$-tree nature of the resulting space. Actually, the $V$-tree ${\mathcal{X}}_{V}({\mathcal{V}})$ will be natuarally a space with peripherals, and the description of a family of peripherals will be a part of its definition.
Similarly as in Section 6, we consider a standard family ${\mathcal{N}}_{\Gamma}$ of normal neighbourhoods in $\Gamma$, as described in Example 2.2, and the space with peripherals $\Gamma^{\circ}$, called the punctured graph $\Gamma$, described in the same example. Consider also the set $\Omega^{V}_{\Gamma}$ of $V$-peripherals of $\Gamma^{\circ}$, as described right after Definition 6.1.
@@
    for each black vertex $t$ of $T$ the constituent space with peripherals at $t$ is equal to the pair $(\Gamma^{\circ},\Omega^{V}_{\Gamma})$; this means in particular that the vertices in $N_{t}$ (i.e. the vertices adjacent to $t$) are in fixed bijective correspondence with the family $\Omega_{\Gamma}^{V}$;
@@
    for each white vertex $u$ of $T$, denote by $e,e^{\prime}$ the two edges of $T$ adjacent to $u$, and by $P_{e},P_{e^{\prime}}$ the $V$-peripherals of $\Gamma_{b(e)}^{\circ}$ and $\Gamma_{b(e^{\prime})}^{\circ}$ associated to $u$ via the bijective correspondences mentioned above in condition (v1); then, denoting by $v,v^{\prime}$ the vertices of $\Gamma$ corresponding to the $V$-peripherals $P_{e},P_{e^{\prime}}$, respectively, we have $a(v)=v^{\prime}$; moreover, the set $\Sigma_{u}$ and the bijections $\varphi_{e}:\Sigma_{u}\to P_{e},\varphi_{e^{\prime}}:\Sigma_{u}\to P_{e^{\prime}}$ are given so that the composition $\varphi_{e^{\prime}}\circ\varphi_{e}^{-1}:P_{e}\to P_{e^{\prime}}$ coincides with the map $\alpha_{v}$, where we refer to the natural identifications $P_{e}={\rm Lk}(v,\Gamma)$ and $P_{e^{\prime}}={\rm Lk}(v^{\prime},\Gamma)$.
We skip a straightforward justification of the fact that a tree system satisfying the above requirements exists and is unique up to an isomorphism of tree systems (the argument is similar to that presented in Section 6 for the tree system $\Theta[{\mathcal{R}}]$; see the proof of Lemma 6.3).
Now, let ${\mathcal{X}}_{V}({\mathcal{V}}):=\lim\Theta[{\mathcal{V}}]$. We describe the family $\Omega_{V}({\mathcal{V}})$ of natural peripheral subsets for ${\mathcal{X}}_{V}({\mathcal{V}})$. Let $P$ be any $E$-peripheral in any constituent space $\Gamma^{\circ}_{t}$, $t\in V^{b}_{T}$. Clearly, $P$ is then naturally a subset of the limit $\lim\Theta[{\mathcal{V}}]$. Moreover, for any two distinct $E$-peripherals $P,P^{\prime}$ as above, the induced subsets of $\lim\Theta[{\mathcal{V}}]$ are disjoint. It is also not hard to realize (by deducing it e.g. from Fact 2.5 in view of Fact 2.6) that the family $\Omega_{V}({\mathcal{V}})$ of all such subsets $P\subset\lim\Theta[{\mathcal{V}}]$, for all $t\in V^{b}_{T}$, is null, and thus we take it as the family of natural peripherals for ${\mathcal{X}}_{V}({\mathcal{V}})$.
@@
A graph which coincides with a single edge will be called the segment, and we donote this graph by $I$. The punctured segment $I^{\circ}$ is then a space with peripherals in which two of the peripheral sets are singletons, while the others are doubletons. If we delete from the set of peripherals of $I^{\circ}$ the two singleton peripherals, we get a new space with peripherals (with the same underlying space, but smaller family of peripherals) which we call the internally punctured segment, and which we denote $I^{\circ}_{\rm int}$. Alternatively, $I^{\circ}_{\rm int}$ can be described as obtained by deleting from $I$ the interiors of a dense and null family of pairwise disjoint normal neighbourhoods of points lying in the interior of $I$, and by taking the boundaries of those neighbourhoods as peripherals.
Observe that there are exactly two essentially distinct connecting $V$-systems ${\mathcal{V}}=(\Gamma,{\mathbf{a}})$ with $\Gamma=I$: the first one with the involution $a:V_{I}\to V_{I}$ equal to the identity of the vertex set $V_{I}$, and the second one with $a$ equal to the transposition of the doubleton $V_{I}$. In both cases the resulting $V$-tree ${\mathcal{X}}_{V}({\mathcal{V}})$ will be called a $V$-tree of segments. Moreover, we obviously have the following.
@@
We now pass to describing some natural decomposition of any $V$-tree of graphs ${\mathcal{X}}_{V}({\mathcal{V}})$ into convenient pieces. These pieces will be called the core and the arms of ${\mathcal{X}}_{V}({\mathcal{V}})$, and they will have a form of spaces with peripherals. We will show (in Lemma 7.12) that decomposition into these pieces has a tree-like nature, i.e., that the space ${\mathcal{X}}_{V}({\mathcal{V}})$ is canonically homeomorphic to a certain tree of these pieces. This new expression of ${\mathcal{X}}_{V}({\mathcal{V}})$ as a tree of spaces does not coincide with the original expression of this space as a $V$-tree of graphs, and this reinterpretation of the spaces of the form ${\mathcal{X}}_{V}({\mathcal{V}})$ will be one of our tools in the proof of the implications (3$\Rightarrow$(2) in Theorem A and (3$\Rightarrow$(2) in Theorem B.
Before describing the pieces and the decomposition, we need to introduce more terminology. A punctured edge of a punctured graph $\Gamma^{\circ}$ is the intersection of $\Gamma^{\circ}$ (viewed as subspace of $\Gamma$) with any edge $e$ of $\Gamma$, and we denote such a punctured edge by $e^{\circ}$. Obviously, since we assume that $\Gamma$ contains no loop edges, each $e^{\circ}$ naturally has the structure of a space with peripherals with which it is homeomorphic to the punctured segment $I^{\circ}$. The endpoints of a punctured edge $e^{\circ}$ are the two points of $e^{\circ}$ which correspond to the two singleton peripherals of this space. The set of all endpoints in all punctured edges of $\Gamma^{\circ}$ is clearly in a natural bijective correspondence with the union of all vertex links of the graph $\Gamma$, and we will identify each such endpoint with the corresponding point in the appropriate vertex link. Now, given a $V$-tree system $\Theta[{\mathcal{V}}]=(T,\{\Gamma_{t}^{\circ}\},\{\Sigma_{u}\},\{\varphi_{e}\})$, we say that punctured edges $e_{1}^{\circ}$ of $\Gamma_{t_{1}}^{\circ}$ and $e_{2}^{\circ}$ of $\Gamma_{t_{2}}^{\circ}$ are $\Theta$-adjacent if some endpoints $x_{1},x_{2}$ in $e_{1}^{\circ},e_{2}^{\circ}$ get identified in the limit space ${\mathcal{X}}_{V}({\mathcal{V}})$, i.e., if $t_{1},t_{2}$ are distinct black vartices of $T$ with common adjacent white vertex, which we denote $u$, and if we have $x_{i}\in\varphi_{[t_{i},u]}(\Sigma_{u})$, for $i=1,2$, and $\varphi_{[t_{1},u]}^{-1}(x_{1})=\varphi_{[t_{2},u]}^{-1}(x_{2})$. We then say that $e_{1}^{\circ}$ and $e_{2}^{\circ}$ are $\Theta$-adjacent at their endpoints $x_{1},x_{2}$. Note that if $t_{1},t_{2},u$ are as above, and if $e_{1}^{\circ}$ is a punctured edge of $\Gamma_{t_{1}}^{\circ}$ having an endpoint $x_{1}\in\varphi_{[t_{1},u]}(\Sigma_{u})$, then there is precisely one punctured edge $e_{2}^{\circ}$ in $\Gamma_{t_{2}}^{\circ}$ which is $\Theta$-adjacent to $e_{1}^{\circ}$.
@@
A line in a $V$-tree system $\Theta[{\mathcal{V}}]$ is a bi-infinite sequence $(e_{n}^{\circ})_{n\in\mathbb{Z}}$ of punctured edges from the punctured graphs $\Gamma_{t}^{\circ}:t\in V^{b}_{T}$ such that for each $n$ the edge $e_{n}^{\circ}$ is $\Theta$-adjacent to $e_{n-1}^{\circ}$ at one of its endpoints, and it is $\Theta$-adjacent to $e_{n+1}^{\circ}$ at its other endpoint. We do not distinguish sequences as above which differ by a shift or by inversion of the indices (i.e. we view two such sequences as representing the same line).
Obviously, since $\Gamma$ has no loop edges, any punctured edge in any constituent graph $\Gamma_{t}^{\circ}$ belongs to a unique line in $\Theta[{\mathcal{V}}]$. It is also not hard to observe that in each $V$-tree system $\Theta[{\mathcal{V}}]$ for which $\Gamma$ has more than two vertices, there are many (actually countably infinitely many) lines. We denote the set of all lines in $\Theta[{\mathcal{V}}]$ by ${\mathcal{L}}({\mathcal{V}})$. Since each term $e_{n}^{\circ}$ of any line $L$ is a punctured edge in some copy $\Gamma_{t}$ of the graph $\Gamma$, we will often refer to it as an appearance in $L$ of the corresponding (unpunctured) edge of $\Gamma$.
Let $L=(e^{\circ}_{n})_{n\in\mathbb{Z}}$ be a line in $\Theta[{\mathcal{V}}]$, and for each $n$ denote by $t_{n}$ this black vertex of $T$ for which $e_{n}^{\circ}$ is a punctured edge of $\Gamma_{t_{n}}^{\circ}$, and by $u_{n}$ the unique white vertex of $T$ adjacent to both $t_{n-1}$ and $t_{n}$. The bi-infinite sequence
@@
is then a (bi-infinite) combinatorial geodesic in $T$, which we denote by $\gamma_{L}$ and call the geodesic induced by $L$. (We view such geodesics as subsets in the vertex set $V_{T}$ rather than as sequences, so that a shift of the indices or reversal of the order of terms yields the same geodesic.) We make a record of the following properties of lines and their induced geodesics, whose straightforward proofs we omit. (These proofs are based on the easily seen observation that lines, viewed as sequences of appearances of edges from $\Gamma$, are periodic.)
@@
An arm in a $V$-tree of graphs ${\mathcal{X}}_{V}({\mathcal{V}})$ is its subspace $A_{L}$ of the following form. Given any line $L\in{\mathcal{L}}({\mathcal{V}})$, $L=(e_{n}^{\circ})_{n\in Z}$, take as $A_{L}$ the union of the set $\partial\gamma_{L}\subset\partial T$ (consisting of the two ends of the geodesic $\gamma_{L}$) and the image in #$\#\Theta[{\mathcal{V}}]$ of the union $\bigcup_{n\in Z}e_{n}^{\circ}$. We view $A_{L}$ as a space with peripherals, where the family of peripherals consists of only one peripheral set, namely $\partial\gamma_{L}$. We distinguish also in $A_{L}$ some other subsets, namely the ones corresponding to all doubleton peripherals in all punctured edges $e_{n}^{\circ}$ constituting $A_{L}$; we call these subsets the supplementary peripherals of $A_{L}$, in contrast with the above mentioned subset $\partial\gamma_{L}$, which we call the basic peripheral of $A_{L}$.
For each arm $A_{L}$ we consider the subspace topology induced from ${\mathcal{X}}_{V}({\mathcal{V}})$. It is not hard to observe by referring to Fact 2.5 that, when equipped with this topology, any arm $A_{L}$ naturally gets the form of a $V$-tree of segments. In particular, by Fact 7.3, we obtain the following.
@@
The core of ${\mathcal{X}}_{V}({\mathcal{V}})=\lim\Theta[{\mathcal{V}}]$ is the subspace $C({\mathcal{V}})$ corresponding to the subset $\partial T\subset\lim\Theta[{\mathcal{V}}]$. It is equipped (as part of its structure) with the family of doubleton subsets $\Omega^{c}_{\mathcal{V}}:=\{\partial\gamma_{L}:L\in{\mathcal{L}}({\mathcal{V}})\}$ (with which it is a space with peripherals, see Remark 7.9.2 below).
@@
    The family $\Omega^{c}_{\mathcal{V}}$ of subsets of the core is null. To see this, it is sufficient to note that the distances in $T$ from a fixed base point to the geodesics $\gamma_{L}$ diverge to $\infty$. This, in turn, easily follows from the fact that only finitely many geodesics of the form $\gamma_{L}$ passes through any vertex of $T$, and the fact that $T$ is locally finite. The family $\Omega^{c}_{\mathcal{V}}$ consists also of sets that are pairwise disjoint, see Lemma 7.5(1).
@@
We now generalize V-trees of segments to V-trees of $\theta$-graphs. We start with describing the $\theta$-graphs—a family of graphs which appear explicitly in the statement of Theorem B, and which also appear implicitly in Theorem A, since the regular trees of graphs mentioned in conditions (2) and (3) of this result necessarily involve certain $\theta$-graphs (compare Theorem 10.1 and the description of the graphical connecting system ${\mathcal{R}}_{G}$ in Section 10).
@@
Generalizing the exposition of Example 7.2, we define the internally punctured $\theta$-graph $(\Xi_{k})^{\circ}_{{\rm int}}$ as the space with peripherals obtained from $\Xi_{k}$ by deleting the interiors of any dense family of pairwise disjoint normal neighbourhoods of non-vertex points of $\Xi_{k}$, and by taking the boundaries of those neighbourhoods as peripherals. Such a space is obviously unique up to a homeomorphism respecting peripherals, its peripherals are all doubletons, and vertices of $\Xi_{k}$ belong to this space, and are not contained in any peripheral.
Observe that any connecting $V$-system ${\mathcal{V}}=(\Xi_{k},{\mathbf{a}})$ of $\theta$-graphs has precisely $k$ lines $L_{1},\dots,L_{k}$. Moreover, the arms $A_{1},\dots,A_{k}$ of the corresponding $V$-tree of $\theta$-graphs ${\mathcal{X}}_{V}({\mathcal{V}})$ have a common pair of endpoints, and this pair coincides with the core $C({\mathcal{V}})$ of this $V$-tree. The set of peripherals of ${\mathcal{X}}_{V}({\mathcal{V}})$ coincides with the union of the sets of supplementary peripherals of the arms $A_{i}$. Thus, as a space with peripherals, ${\mathcal{X}}_{V}({\mathcal{V}})$ is homeomorphic to the internally punctured $\theta$-graph $(\Xi_{k})^{\circ}_{\rm int}$. The core $C({\mathcal{V}})$ of this $V$-tree coincides then with the vertex set of $\Xi_{k}$ (viewed as the subset of $(\Xi_{k})^{\circ}_{\rm int}$). The only peripheral of this core coincides with the whole of the core.
@@
As the constituent spaces for $\Psi[{\mathcal{V}}]$ we take $K_{t_{c}}=C({\mathcal{V}})$ and $K_{t_{A}}=A$. As the peripheral spaces for $\Psi[{\mathcal{V}}]$ we take $\Sigma_{u_{P}}=P$, and as the maps $\varphi_{[u,t]}:\Sigma_{u}\to K_{t}$ we take the natural inclusions of the forms $P\subset C[{\mathcal{V}}]$ and $\partial\gamma_{L}\subset A_{L}$.
@@
The inclusions of the form $A_{L}\subset{\mathcal{X}}_{V}({\mathcal{V}})$, for all $L\in{\mathcal{L}}({\mathcal{V}})$, and $C[{\mathcal{V}}]\subset{\mathcal{X}}_{V}({\mathcal{V}})$ induce a well defined map $i_{\mathcal{V}}:\lim\Psi[{\mathcal{V}}]\to{\mathcal{X}}_{V}({\mathcal{V}})$, and this map is a homeomorphism.